# Efficiently Editing Mixture-of-Experts Models with Compressed Experts

### v1

@@ 5 Conclusions
In this work, we introduce compressed expert, a lightweight and efficient approach to reducing the number of activated experts in MoE models while maintaining strong performance. By replacing auxiliary experts with compact expert representations, our method significantly reduces computational overhead while preserving model capacity. Extensive experiments on Phi-MoE and OLMoE across various tasks demonstrate that compressed experts reduce active parameters by over 30% and cut inference costs by 20% while retaining over 90% of the performance of full-expert configurations. Beyond reducing inference costs, our findings suggest broader implications for scaling MoE models efficiently. Compressed experts offer a promising direction for optimizing sparse activation in MoE architectures.
@@
Our approach uses a fixed compression ratio, which treats half of the activated experts as main experts, while the other half are auxiliary experts. As shown in Section 3.2, this configuration may not be optimal for all tasks, especially for models with more than 2 activated experts, which enables more possibilities of main and auxiliary expert distribution. While halving provides a strong efficiency-performance trade-off from our empirical validation in Section 2.3, some tasks may require a more adaptive compression strategy. To tackle this challenge, future works can explore dynamically determining the degree of expert reduction based on task complexity or input characteristics, and address the additional computational overhead incurred by those approaches.
@@ Appendix A Active Parameter calculation
Here, we provide a similar calculation of active parameter reduction for OLMoE as in Section 2.4. In OLMoE, each MoE layer contains 3 weight matrices of size $1024\times 2048$. The total non-MoE parameters are approximately 475M. In the original configuration with 8 activated experts across 16 MoE layers, the number of parameters in MoE layers is
@@
Appendix B Dataset Details
@@
Appendix C Potential Risks

### v2

@@ 5 Conclusions
We introduce compressed expert, a lightweight and efficient approach to reducing the number of activated experts in MoE models while maintaining strong performance. By replacing auxiliary experts with compact expert representations, our method significantly reduces computational overhead while preserving model capacity. Extensive experiments on Phi-MoE and OLMoE across various tasks demonstrate that compressed experts reduce active parameters by over 30% and cut inference costs by 20% while retaining over 90% of the performance of full-expert configurations. Beyond reducing inference costs, our findings suggest broader implications for scaling MoE models efficiently. Compressed experts offer a promising direction for optimizing sparse activation in MoE architectures.
@@
Our approach uses a fixed compression ratio, which treats half of the activated experts as main experts, while the other half are auxiliary experts. As shown in Section˜3.2, this configuration may not be optimal for all tasks, especially for models with more than 2 activated experts, which enables more possibilities of main and auxiliary expert distribution. While halving provides a strong efficiency-performance trade-off from our empirical validation in Section˜2.3, some tasks may require a more adaptive compression strategy. To tackle this challenge, future works can explore dynamically determining the degree of expert reduction based on task complexity or input characteristics, and address the additional computational overhead incurred by those approaches.
@@ Appendix A Active Parameter Calculation
Here, we provide a similar calculation of active parameter reduction for OLMoE as in Section˜2.4. In OLMoE, each MoE layer contains 3 weight matrices of size $1024\times 2048$. The total non-MoE parameters are approximately 475M. In the original configuration with 8 activated experts across 16 MoE layers, the number of parameters in MoE layers is
@@
Appendix B Performance-Latency Trade-off
Figure 5: The performance of OLMoE versus the inference latency, each point representing a different expert configuration. The Top-4 w/ CE configuration performs closely to the Top-8 configuration while achieving low inference latency close to Top-4.
We produce a similar plot as Figure˜4 for OLMoE. Similar to our observations with Phi-MoE, incorporating compressed experts into the Top-4 configuration strikes a favorable balance between efficiency and performance. Specifically, It only adds a minimal overhead on the inference latency compared with top-4, but noticeably closes the performance gap to top-8. This consistent trend across different models further validates the effectiveness of our approach.
Appendix C Alternative Compressed Experts Construction
During the development of our compressed expert design, we have explored alternatives such as autoencoder reductions or LoRA. However, we find that those approaches do not noticeably decrease the inference time. While these methods reduce the parameter count, they still require the execution of full forward passes through MLPs layers, similar to the full experts, thereby retaining much of the original computational overhead. For instance, we compare the inference latency of a LoRA-based implementation with rank 16 for Phi-MoE against our compressed expert approach in Table˜7. As shown, LoRA still introduces a noticeable increase in inference time.
Top-1 Top-1 W/CE Top-1-W/LoRA Top-2
Inference latency (s) 4.01 4.35 5.42 5.59
Table 7: Inference latency for alternative CE construction.
In contrast, our element-wise multiplication approach requires only a simple, low-cost operation to incorporate compressed experts into the model. Empirically, we found that this approach retains sufficient expressive power to capture the critical information from auxiliary experts while substantially lowering inference costs. We will include this discussion in our final manuscript.
Appendix D Training details
Appendix E Dataset Details
@@
Appendix F Potential Risks